# Data Contamination Issues in Brain-to-Text Decoding

### 18 Dec 2023 v1

@@ Abstract
Decoding non-invasive cognitive signals to natural language has long been the goal of building practical brain-computer interfaces (BCIs). Recent major milestones have successfully decoded cognitive signals like functional Magnetic Resonance Imaging (fMRI) and electroencephalogram (EEG) into text under open vocabulary setting. However, how to split the datasets for training, validating, and testing in cognitive signal decoding task still remains controversial. In this paper, we conduct systematic analysis on current dataset splitting methods and find the existence of data contamination largely exaggerates model performance. Specifically, first we find the leakage of test subjects’ cognitive signals corrupts the training of a robust encoder. Second, we prove the leakage of text stimuli causes the auto-regressive decoder to memorize information in test set. The decoder generates highly accurate text not because it truly understands cognitive signals. To eliminate the influence of data contamination and fairly evaluate different models’ generalization ability, we propose a new splitting method for different types of cognitive datasets (e.g. fMRI, EEG). We also test the performance of SOTA Brain-to-Text decoding models under the proposed dataset splitting paradigm as baselines for further research.
@@ 1 Introduction
Brain-computer interface (BCI) builds connections between cognitive signals of human brain and external devices (e.g. computer). It has been widely researched in the field of neuroscience and has gained remarkable success like repairing damaged sight or restoring movement of disabled people (Polikov et al., 2005; Hochberg et al., 2012; Bouton et al., 2016). However, it’s still challenging in restoring subjects’ cognitive signals when they read or hear text stimuli back to corresponding natural language, especially when it comes to non-invasive cognitive signals like functional Magnetic Resonance Imaging (fMRI) or electroencephalogram (EEG), which are noisy and of low resolution (Mridha et al., 2021).
@@
Figure 1: General frameworks of decoding cognitive signals to text and possible ways of data contamination.
@@
Previous work (Herff et al., 2015; Anumanchipalli et al., 2019; Zou et al., 2021; Moses et al., 2021; Défossez et al., 2023) mainly focused on word-level cognitive signal decoding with restricted vocabulary size (hundreds of words). Although these models achieve high accuracy, they struggle to generate unseen words, thus leading to poor generalization. Recent methods (Wang and Ji, 2022; Xi et al., 2023) leveraged powerful pre-trained language models like BART (Lewis et al., 2020) in decoding cognitive signals to text and achieved open vocabulary sentence-level generation. These methods typically view the cognitive signal decoding task as machine translation (Sutskever et al., 2014; Bahdanau et al., 2015; Makin et al., 2020) and adopt an encoder-decoder framework, where the encoder is responsible for converting cognitive signals into low-dimensional representations and the decoder learns to map the representations to natural language. Specifically, as shown in Figure1, the encoder usually consists of a spatial and time series feature extractor. It can be trained either in an end-to-end manner with decoder (Figure 1 (a)) or first pre-trained through a signal reconstruction task and then applied in decoder training (Figure 1 (b)). Despite recent success in model design, it still remains controversial in how to split the dataset for training, validating, and testing (Xi et al., 2023). Addressing this issue is urgent and meaningful, because fair evaluations and comparisons of different models are impossible without a widely recognized dataset splitting paradigm.
@@
Datasets are usually formatted as signal-sentence pairs in decoding cognitive signals to text task. In most cases, each sentence belongs to a certain task, so signal-sentence pairs can be further divided into signal-task and task-sentence pairs. Current dataset splitting methods (Wang and Ji, 2022; Xi et al., 2023) can be summarized into five categories: (1) split by subjects (2) split by tasks (3) split by randomly picking signal frames (4) split by randomly picking signal frames under certain task (5) split by randomly picking consecutive signal frames under certain task. However, all these methods suffer from data contamination on encoder side, decoder side, or both. As to encoder, if subjects’ cognitive signals in test set are mixed to training set, encoder will become overfitted and fail to well represent unseen subjects’ cognitive signals. As to decoder, situation gets worse if text stimuli is leaked. Since the decoder generates token by token in an auto-regressive manner, data contamination will cause the decoder to memorize seen paragraphs and probability distribution, which means given the first few tokens the decoder can predict next tokens without encoded cognitive signal representations.
@@
To address the problems, we propose a new dataset splitting method that eradicates data contamination from both encoder and decoder sides. We focus on fMRI and EEG signals in experiments, although the proposed splitting method could be applied to any cognitive signals satisfying the above-mentioned format. In our method of dataset partition, dataset is split according to subject-stimuli pairs with the following rules: (1) Cognitive signals collected from subjects in validation and test set won’t appear in training set, which means the trained encoder can’t access any brain information belonging to the subjects in test set. (2) Text stimuli in validation and test set won’t appear in training set. The decoder learns the mapping between cognitive signal representation and token embedding instead of memorizing seen segment.
@@
  • We investigate current dataset splitting methods and analyze the effect on existing frameworks in decoding cognitive signals to text.
  • We prove data contamination exists in current dataset splitting methods through analysis and experiments, which seriously exaggerates model performance.
  • We propose the first dataset splitting method without data contamination. We also release a fair benchmark to evaluate different models’ generalization performance for further research in this domain.
@@ 2 Related Work
Cognitive Signals Cognitive signals can be classified into three categories: invasive (microelectrode array), partially invasive (ECoG) and non-invasive (EEG, MEG, fMRI) according to how close electrodes get to brain tissue. Due to the high cost and complexity of invasive and partially invasive methods, it’s hard to apply them in building generic and practical BCIs. Here we mainly focus on non-invasive signals EEG and fMRI. EEG signal is electrogram of the spontaneous electrical activity of the brain. Its frequencies usually range from 1 to 30 Hz, divided into several groups like alpha (4-13 Hz), beta (13-30 Hz), delta (0.5-4 Hz), theta (4-7 Hz). EEG is of high temporal resolution and relatively tolerant of subject movement, but its spatial resolution is low and it can’t display active areas of the brain directly. fMRI measures brain activity by detecting changes of blood flow. Blood flow of a specific region increases when this brain area is in use. The spatial resolution of fMRI is measured by the size of voxel, which is a three-dimensional rectangular cuboid ranging from 3mm to 5mm (Vouloumanos et al., 2001; Noppeney and Price, 2004). Unlike EEG which samples brain signals continuously, fMRI samples according to the smallest time period named TR, usually at second level.
Non-invasive Brain Decoding Previous researches (Herff et al., 2015; Anumanchipalli et al., 2019; Zou et al., 2021; Moses et al., 2021; Défossez et al., 2023) on brain-to-text decoding mainly focused on word-level decoding in a restricted vocabulary with hundreds of words (Panachakel and Ramakrishnan, 2021). These models typically apply recurrent neural network or long short-term memory (Hochreiter and Schmidhuber, 1997) network to build mapping between cognitive signals and words in vocabulary. Despite relatively good accuracy, these methods fail to generalize to unseen words. Some progress (Sun et al., 2019) has been made by expanding word-level decoding to sentence-level through encoder-decoder framework, or use less noisy ECoG data (Burle et al., 2015; Anumanchipalli et al., 2019). However, these models struggle to generate accurate and fluent sentences limited by decoder ability. Wang and Ji (2022) introduced the first open vocabulary EEG-to-text decoding model by leveraging the power of pre-trained language models. Xi et al. (2023) improved the model design and proposed a unified framework for decoding both fMRI and EEG signals.
@@ 3 Methodology
In this section, we will first introduce the definition of decoding cognitive signals to text task and the general description of dataset format. Then we systematically analyze current dataset splitting methods and point out that all existing methods suffer from two kinds of data contamination: cognitive signal leakage and text stimuli leakage. Finally, a new dataset splitting method is proposed to avoid the above-mentioned data contamination.
@@ 3.1 Task Definition
Given the cognitive signal $F_{ij}$ stimulated by $i$-th subject $S_{i}$ hearing or reading certain text $T_{j}$, the goal is to decode $F_{ij}$ back to text $T_{j}^{\prime}$ and make $T_{j}^{\prime}$ as similar as possible to $T_{j}$. The composition of $F_{ij}$ and $T_{j}$ is different as to fMRI and EEG signals. The former samples brain information discretely with a fixed time interval TR, while the latter samples continuously. To fMRI, $L_{window}$ consistent sentence segments $s_{j}$ with fMRI frames $f_{ij}$ are concatenated to form a sample pair $\left\langle F_{ij},T_{j}\right\rangle$. Here $T_{j}=concat(s_{j},\ldots,s_{j+L_{window}-1})$ and $F_{ij}=concat(f_{ij},\ldots,f_{ij+L_{window}-1})$. To EEG, since signals corresponding to a complete sentence are available, we bond sentence (aka text stimuli) $T_{j}$ and EEG signal $F_{ij}$ together to form a sample pair $\left\langle F_{ij},T_{j}\right\rangle$. Under most circumstances, each sentence $T_{j}$ belongs to a certain task $M_{k}$. So the signal-sentence pair $\left\langle F_{ij},T_{j}\right\rangle$ can be further split into $\left\langle F_{ij},M_{k}\right\rangle$ and $\left\langle M_{k},T_{kj}\right\rangle$.
In decoding cognitive signals to text, the ultimate goal of trained BCI models is to generalize to unseen subjects with unseen text stimuli (Huang et al., 2010; Handiru and Prasad, 2016; Gao et al., 2021). If cognitive signal $F_{ij}$ appears in test set $S_{test}$, then any subject $i$’s signals $F_{i}$ should not appear in training set $S_{train}$. Similarly, text stimuli $T_{kj}$ in $S_{test}$ should not appear in $S_{train}$. So the dataset splitting rules can be formally defined as:
@@ 3.2 Dataset Splitting Methods
Current dataset splitting methods can be summarized as five categories according to classifying objective $S_{i},M_{k},T_{kj},F_{ij}$. More specifically, five dataset splitting methods are named as (1) split by subjects (2) split by tasks (3) split by randomly picking signal frames (4) split by randomly picking signal frames under certain task (5) split by randomly picking consecutive signal frames under certain task, corresponding to image (a), (b), (c), (d), (e) in Figure 2. Figure 2 is plotted to vividly display the difference between current dataset splitting methods. For simplicity of expression, we choose four subjects with three tasks each containing four, three, four sentences in Figure 2. The line connecting two symbols indicates they are related to a sample. Take path $S_{1},M_{1},T_{11},F_{11}$ for example, it’s a sample where $S_{1}$ listens to stimuli $T_{11}$ belonging to $M_{1}$ and $S_{1}$’s brain signal is recorded as $F_{11}$. Some symbols are connected with several lines. For example, the four lines between $S_{1}$ and $M_{1}$ correspond to $\left\langle M_{1},T_{11}\right\rangle$, $\left\langle M_{1},T_{12}\right\rangle$, $\left\langle M_{1},T_{13}\right\rangle$, $\left\langle M_{1},T_{14}\right\rangle$ counting from left to right. Similarly, the three lines between $M_{1}$ and $T_{11}$ correspond to $\left\langle S_{1},M_{1}\right\rangle$, $\left\langle S_{2},M_{1}\right\rangle$, $\left\langle S_{3},M_{1}\right\rangle$ respectively. The same rules can be extended to other lines and symbols. The green lines and orange lines in Figure 2 stand for training samples and testing samples. The grey dotted line means the sample is abandoned, which will be introduced in our splitting method. As the splitting of validation set is similar to test set, we only consider training set and test set in this section.
@@
for training set. Method (c), (d), and (e) all split the dataset according to cognitive signal frames, which can be described as
$S_{train}=\{\left\langle F_{ij},T_{kj}\right\rangle|F_{ij}\notin S_{val},S_{test}\}$ (6)
for training set. However, there are slight differences between these methods. Method (c) views all the cognitive signal frames in dataset as a whole and splits according to the default proportion (e.g. 8:1:1). Method (d) views signal frames under certain task $M_{k}$ as a whole and splits proportionally. Then union all training sets under different tasks to form a complete set for training. Method (e) is similar to method (d). They both first split training, validation, and test set under certain task proportionally and then union them. The difference lies in that method (d) randomly picks signal frames while method (e) consecutively picks signal frames.
@@
Table 1: Results of Cognitive Signal Leakage Rate (CSLR) and Text Stimuli Leakage Rate (TSLR)
The goal of decoding cognitive signals to text models is to generalize to unseen subjects with unseen text stimuli, which means both subject’s brain information and received stimuli are new to the trained model. In this sense, we define two kinds of data contamination: cognitive signal leakage and text stimuli leakage. The former will harm encoder training, while the latter will pollute the decoder. The data contamination condition of different methods is reflected in Figure 2. If lines connecting two symbols are of different colours, it means data in test set leaks into training set. Lines between $S_{i}$ and $M_{k}$ indicate cognitive signal leakage condition and lines between $T_{kj}$ and $M_{k}$ indicate text stimuli leakage condition. As shown in the figure, cognitive signal leakage doesn’t exist in method (a), but method (a) suffers from text stimuli leakage. The situation of method (b) is opposite to that of method (a). There’s no text stimuli leakage but cognitive signal leakage. Method (c) and method (d) are similar. They suffer from both cognitive signal leakage and text stimuli leakage. Method (e) is in the same situation as method (b). The composition of samples differs between fMRI signals and EEG signals, so the dataset splitting methods for two cognitive signals are different too. Since fMRI signals need to be sampled continuously with a certain length $L_{window}$, the path of a sample shown in Figure 2 is actually the first part of one fMRI sample, with $L_{window}-1$ continuous part following. In this sense, method (c), (d), and (e) which seem the same for EEG are different for fMRI. Besides, in method (e) there still exists slight text stimuli leakage in the overlap between training samples and test samples for fMRI signal.

### 26 Dec 2023 v2

@@ Abstract
Decoding non-invasive cognitive signals to natural language has long been the goal of building practical brain-computer interfaces (BCIs). Recent major milestones have successfully decoded cognitive signals like functional Magnetic Resonance Imaging (fMRI) and electroencephalogram (EEG) into text under open vocabulary setting. However, how to split the datasets for training, validating, and testing in brain-to-text decoding still remains controversial. In this paper, we conduct systematic analysis on current dataset splitting methods and find data contamination largely exaggerates model performance. Specifically, first we find the leakage of test subjects’ cognitive signals corrupts the training of a robust encoder. Second, we prove the leakage of text stimuli causes the auto-regressive decoder to memorize seen information in test set. To eliminate the influence of data contamination and fairly evaluate different models’ generalization ability, we propose a new splitting method for different types of cognitive dataset (e.g. fMRI, EEG). We also test the performance of SOTA brain-to-text decoding models under the proposed dataset splitting paradigm as baselines for further research.
@@ 1 Introduction
Brain-computer interface (BCI) builds connections between human brain and external devices (e.g. computer). It has been widely researched in the field of neuroscience and has gained remarkable success like repairing damaged sight or restoring movement of disabled people (Polikov et al., 2005; Hochberg et al., 2012; Bouton et al., 2016). However, it’s still challenging in decoding cognitive signals to corresponding natural language when subjects read or hear text stimuli, especially for non-invasive cognitive signals like functional Magnetic Resonance Imaging (fMRI) or electroencephalogram (EEG) that are noisy and of low resolution (Mridha et al., 2021).
@@
Figure 1: General frameworks of brain-to-text decoding and possible situations of data contamination.
@@
Recent methods (Makin et al., 2020; Wang and Ji, 2022; Xi et al., 2023; Tang et al., 2023) typically viewed brain-to-text decoding as machine translation (Sutskever et al., 2014; Bahdanau et al., 2015) and adopted an encoder-decoder framework, where the encoder is responsible for converting cognitive signals into low-dimensional representations and the decoder learns to map the representations to natural language. Specifically, as shown in Figure1, the encoder usually consists of a spatial and time series feature extractor. It can be trained either in an end-to-end manner with decoder (Figure 1 a) or first pre-trained through a signal reconstruction task and then applied in decoder training (Figure 1 b). Despite recent success in model design, it still remains controversial in how to split the dataset for training, validating, and testing (Xi et al., 2023). Addressing this issue is urgent and meaningful, as fair evaluation of models is impossible without a widely recognized dataset splitting paradigm.
@@
A cognitive dataset is usually formatted in signal-sentence pair. In most cases for brain-to-text decoding task, each sentence belongs to a certain task, so signal-sentence pair can be further divided into signal-task and task-sentence pair. Current dataset splitting methods (Wang and Ji, 2022; Xi et al., 2023) can be summarized into five categories: (1) split by subjects, (2) split by tasks, (3) split by randomly picking signal frames, (4) split by randomly picking signal frames under certain task, (5) split by randomly picking consecutive signal frames under certain task. However, all these methods suffer from data contamination on encoder side, decoder side, or both. As shown in Figure 1, to encoder if subjects’ cognitive signals in test set are mixed into training set, encoder will become overfitted and fail to well represent unseen subjects’ cognitive signals. As to decoder, situation gets worse if text stimuli are leaked. Since the decoder generates token by token in an auto-regressive manner, data contamination will cause the decoder to memorize seen paragraphs and probability distribution, which means given the first few tokens the decoder is able to predict next token regardless of encoded cognitive signal representations.
@@
To address the above-mentioned problems, we propose a new dataset splitting method that eradicates data contamination from both encoder and decoder sides. We focus on fMRI and EEG signals in experiments, although the proposed splitting method could be applied to any cognitive signals satisfying the given format. In our method, the dataset is split according to subject-stimuli pairs with the following rules: (1) Cognitive signals collected from subjects in validation set and test set won’t appear in training set, which means the trained encoder can’t get access to any brain information belonging to subjects in test set. (2) Text stimuli in validation set and test set won’t appear in training set. The decoder learns the mapping between cognitive signal representation and token embedding instead of memorizing seen text.
@@
  • We investigate current dataset splitting methods and analyze their influence on popular frameworks in brain-to-text decoding.
  • We prove the existence of data contamination in current dataset splitting methods through analysis and experiments, which seriously exaggerates model performance.
  • We propose the first splitting method without data contamination on public cognitive datasets. We also release a fair benchmark to evaluate different models’ generalization performance for further research in this domain.
@@ 2 Related Work
Cognitive Signal Cognitive signals can be classified into three categories: invasive, partially invasive, and non-invasive according to how close electrodes get to brain tissue. Due to the high cost and complexity of invasive and partially invasive methods, it’s hard to apply them in building generic and practical BCIs. In this paper, we mainly focus on non-invasive signals EEG and fMRI. EEG signal is electrogram of the spontaneous electrical activity of the brain. Its frequencies usually range from 1 to 30 Hz, divided into several groups like alpha (4-13 Hz), beta (13-30 Hz), delta (0.5-4 Hz), theta (4-7 Hz). EEG is of high temporal resolution and relatively tolerant of subject movement, but its spatial resolution is low and it can’t display active areas of the brain directly. fMRI measures brain activity by detecting changes of blood flow. Blood flow of a specific region increases when this brain area is in use. The spatial resolution of fMRI is measured by the size of voxel, which is a three-dimensional rectangular cuboid ranging from 3mm to 5mm (Vouloumanos et al., 2001; Noppeney and Price, 2004). Unlike EEG which samples brain signals continuously, fMRI samples according to the smallest time period named TR, usually at second level.
Brain-to-text Decoding Previous research on brain-to-text decoding (Herff et al., 2015; Anumanchipalli et al., 2019; Zou et al., 2021; Moses et al., 2021; Défossez et al., 2023) mainly focused on word-level decoding in a restricted vocabulary with hundreds of words (Panachakel and Ramakrishnan, 2021). These models typically apply recurrent neural network or long short-term memory (Hochreiter and Schmidhuber, 1997) network to build mapping between cognitive signals and words in vocabulary. Despite relatively good accuracy, these methods fail to generalize to unseen words. Some progress (Sun et al., 2019) has been made by expanding word-level decoding to sentence-level through encoder-decoder framework, or use less noisy ECoG data (Burle et al., 2015; Anumanchipalli et al., 2019). However, these models struggle to generate accurate and fluent sentences limited by decoder ability. Wang and Ji (2022) introduced the first open vocabulary EEG-to-text decoding model by leveraging the power of pre-trained language models. Xi et al. (2023) improved the model design and proposed a unified framework for decoding both fMRI and EEG signals.
@@ 3 Methodology
In this section, we will first introduce the definition of decoding cognitive signals to text and the general description of dataset format. Then we systematically analyze current dataset splitting methods and point out that all existing methods suffer from two kinds of data contamination: cognitive signal leakage and text stimuli leakage. Finally, a new dataset splitting method is proposed to avoid the above-mentioned two kinds of data contamination.
@@ 3.1 Task Definition
Given the cognitive signal $F_{ij}$ stimulated by $i$-th subject $S_{i}$ hearing or reading certain text $T_{j}$, brain-to-text decoding aims to decode $F_{ij}$ back to text $T_{j}^{\prime}$ and make $T_{j}^{\prime}$ as similar as possible to $T_{j}$. The composition of $F_{ij}$ and $T_{j}$ is different as to fMRI and EEG. The former samples brain information discretely with a fixed time interval TR, while the latter samples continuously. To fMRI, $L$ consistent sentence segments $s_{j}$ with corresponding fMRI frames $f_{ij}$ are concatenated to form a sample pair $\left\langle F_{ij},T_{j}\right\rangle$. Here $T_{j}=concat(s_{j},s_{j+1},\ldots,s_{j+L-1})$ and $F_{ij}=concat(f_{ij},f_{i,j+1},\ldots,f_{i,j+L-1})$. To EEG, since signals corresponding to a complete sentence are available, we bond sentence (aka text stimuli) $T_{j}$ and EEG signal $F_{ij}$ together to form a sample pair $\left\langle F_{ij},T_{j}\right\rangle$. Under most circumstances, each sentence $T_{j}$ belongs to one certain task $M_{k}$. So the signal-sentence pair $\left\langle F_{ij},T_{j}\right\rangle$ can be further split into $\left\langle F_{ij},M_{k}\right\rangle$ and $\left\langle M_{k},T_{kj}\right\rangle$.
Figure 3: The process of our proposed dataset splitting method.
In brain-to-text decoding, the ultimate goal of trained BCI models is to generalize to unseen subjects with unseen text stimuli (Huang et al., 2010; Handiru and Prasad, 2016; Gao et al., 2021). As a result, if cognitive signal $F_{ij}$ appears in test set $S_{test}$, then any subject $i$’s signal $F_{i}$ should not appear in training set $S_{train}$. Similarly, text stimuli $T_{kj}$ in $S_{test}$ should not appear in $S_{train}$. So the dataset splitting rules can be formally defined as:
@@ 3.2 Dataset Splitting Methods
Current dataset splitting methods can be summarized as five categories according to classifying objectives $S_{i},M_{k},T_{kj},F_{ij}$. More specifically, the five dataset splitting methods are characterised as (1) split by subjects, (2) split by tasks, (3) split by randomly picking signal frames, (4) split by randomly picking signal frames under certain task, (5) split by randomly picking consecutive signal frames under certain task, corresponding to image (a), (b), (c), (d), (e) in Figure 2. Figure 2 vividly displays the differences between current dataset splitting methods. For simplicity of expression, we choose four subjects with three tasks each containing four, three, four sentences. The line connecting two symbols indicates they are related to one sample. Take path $S_{1},M_{1},T_{11},F_{11}$ for example, it’s one sample where subject $S_{1}$ listens to text stimuli $T_{11}$ belonging to task $M_{1}$ and $S_{1}$’s corresponding brain signal is recorded as $F_{11}$. Some symbols are connected with several lines. For example, the four lines between $S_{1}$ and $M_{1}$ correspond to $\left\langle M_{1},T_{11}\right\rangle$, $\left\langle M_{1},T_{12}\right\rangle$, $\left\langle M_{1},T_{13}\right\rangle$, $\left\langle M_{1},T_{14}\right\rangle$ counting from left to right. Similarly, the three lines between $M_{1}$ and $T_{11}$ correspond to $\left\langle S_{1},M_{1}\right\rangle$, $\left\langle S_{2},M_{1}\right\rangle$, $\left\langle S_{3},M_{1}\right\rangle$ respectively. The same rules can be extended to other lines and symbols. The green lines and orange lines stand for training samples and testing samples. The grey dotted line means the sample is abandoned, which will be introduced in our dataset splitting method. As the splitting of validation set is similar to test set, we only consider training set and test set in this section.
@@
for training set. Method (c), (d), and (e) all split the dataset according to cognitive signal frames
$S_{train}=\{\left\langle F_{ij},T_{kj}\right\rangle|F_{ij}\notin S_{val},S_{test}\}.$ (6)
However, there are slight differences between these methods. Method (c) views all the cognitive signal frames in dataset as a whole and splits according to the default proportion (e.g. 8:1:1). Method (d) views signal frames under certain task $M_{k}$ as a whole and splits proportionally, and then union all training sets under different tasks to form a complete set for training. Method (e) is similar to method (d). They both first split training, validation, and test set under certain task proportionally and then union them. The difference lies in that method (d) randomly picks signal frames while method (e) picks consecutive signal frames.
@@
Table 1: Results of Cognitive Signal Leakage Rate (CSLR) and Text Stimuli Leakage Rate (TSLR).
The goal of brain-to-text decoding models is to generalize to unseen subjects with unseen text stimuli, which means both subject’s brain information and received text stimuli are new to the trained model. In this sense, we define two kinds of data contamination: cognitive signal leakage and text stimuli leakage. The data contamination situation of different methods is reflected in Figure 2. If lines associated with $S_{i}$ or $T_{kj}$ are of different colours, data in test set leaks into training set. Lines between $S_{i}$ and $M_{k}$ indicate cognitive signal leakage situation and lines between $T_{kj}$ and $M_{k}$ indicate text stimuli leakage situation. Remind the composition of samples differs as to fMRI signal and EEG signal, so the dataset splitting methods are different for two cognitive signals too. Since fMRI signals need to be sampled continuously with a certain length $L$, the path of a sample shown in Figure 2 is actually the first part of one fMRI sample, with $L-1$ continuous part following. In this sense, for EEG cognitive signal leakage doesn’t exist in method (a), but method (a) suffers from text stimuli leakage. The situation of method (b) is opposite to that of method (a), where there’s no text stimuli leakage but cognitive signal leakage. Method (c) and method (d) are similar. They suffer from both cognitive signal leakage and text stimuli leakage. Method (e) is in the same situation as method (b). For fMRI, method (c), (d), and (e) which seem the same for EEG are actually different splitting ways. The situation of data contamination for different methods is similar to EEG, except for method (e) there still exists slight text stimuli leakage in the overlap between training samples and test samples.